OpenStudy (anonymous):

Mike wants to make meatloaf. His recipe uses a total of 8 pounds of meat. If he uses a 3 to 1 ratio of beef to pork, how much pork will he use? Enter your answer as a mixed number in simplest terms. @misz.lopez115 @hartnn

OpenStudy (anonymous):

Let us say that x is the total amount of pork and y is the total amount of beef. That gives us two equations: x + y = 8 3x = y We solve this using a system of linear equation. x + y = 8 3x = y <=> x = 8 - y 3x = y <=> x = 8 - y 3(8 - y) = y <=> x = 8 - y 24 - 3y = y <=> x = 8 - y 24 = 4y <=> x = 8 - y y = 6 <=> x = 8 - 6 = 2 y = 6 Your solution set is {(2, 6)}. This means you use 2 pound of pork and 6 pounds of beef. Algebra takes a lot of practice, but once you get to know how things work, there is a huge amount of problems you can solve easily, without trial-and-error solutions. if it's wrong, im sorry

OpenStudy (joshawad):

\[2 \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

So is the final answer what josh said?

OpenStudy (anonymous):

I'm not sure, I got 2 and or 2/6 but again i could be wrong

OpenStudy (anonymous):

If you got 2 2/6 then that simplifies to 2 1/3 right?

OpenStudy (anonymous):

Right, thanks ^.^

OpenStudy (anonymous):

thx